A mother is heterozygous for an over-expressed plastin allele on her X chromosome, thus has longer than normal axons. She is also heterozygous for the SMN1 gene on chromosome 5. Her children’s father is normal for plastin and has normal length axons. He too is heterozygous for the SMN1 gene. Both genes influence spinal muscular atrophy type 1 (SMA1); two abnormal copies of the SMN1 gene are required to cause spinal muscular atrophy, but their effect is prevented by high expression of the plastin gene (Lewis 92).

a. Give the genotypes of the mother and the father; provide a legend for your symbols. For each gene, explain why the abnormal alleles are dominant or recessive.

b. List all the possible genotypes and proportions of their offspring for the plastin gene.

c. List all the possible genotypes and proportions of their offspring for the SMN1 gene.

d. Showing your work, determine the probability that this couple will have a daughter with SMA.

e. Showing your work, determine the probability that this couple will have a son with SMA.
Answer #1

1. Genotype of mother = AaX+Xp

Genotype of father = AaXpY

A is representing dominant and normal allele for  SMN1 gene.

a is representing recessive and abnormal allele for SMN1 gene

Xp  = normal recessive allele of plastin gene on X chromosome

X+ = dominant abnormal allele of over expressed plastin gene on X chromosome.

Abnormal allele for SMN1 gene is recessive because it is expressed only in the homozygous condition.ie., two abnormal copies of SMN1 gene are required to cause muscular dystrophy.

Abnormal (over expressed) plastin allele is dominant as it is expressed in the mother during heterozygous condition.  

b. Cross for plastin gene

Parental X+Xp  × XpY

F1 X+Xp : X+Y : XpXp : XpY ( genotypes)

1 : 1 : 1 : 1 ( proportion)

c. Cross for SMN 1 gene

Parental Aa × Aa

F1 AA : Aa : aa ( possible genotype)

1 : 2 : 1 (proportion)

d. Daughters would have SMA when they are recessive for both the alleles. Because presence of single abnormal SMN1 allele would prevent the effect. So the genotype would be aaXpXp

  probability for daughter with SMA = 1/16

e. Sons would have SMA when they are recessive for plastin gene and homozygous recessive for SMN1 gene. Their genotype would be aaXpY

Probability for son with SMA = 1/16